OpenStudy (anonymous):

What is the half life of -0.02m?

OpenStudy (solomonzelman):

idk

OpenStudy (wolf1728):

psam24 Don't you have a similar question posted on this site?

OpenStudy (anonymous):

yea is that wrong?

OpenStudy (wolf1728):

No - I just figured it is easier to work in one place.

OpenStudy (wolf1728):

For example in the other question you state the time is in days but nothing is stated here.

OpenStudy (wolf1728):

When you say half-life is -.02m does that mean after one day, .02 of the sample has decayed?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

other instructions were more descriptive

OpenStudy (wolf1728):

See attached graphic: (I'm not exactly sure if I understand the problem correctly, but here's my explanation). Based on the data that after 1 day, .02 of the original sample has decayed, then we can say original amount is 100, present amount = 98 and time is one day. half-life = (time)*log(2)/log(orig amt / present amt) half-life = 1 day * 0.3010299957 / log (100/98) half-life = 1 day * 0.3010299957 / log (100/98) half-life = 1 day * 0.3010299957 / log (1.0204081633) half-life = 0.3010299957 / 0.0087739243 half-life = 34.31 days (Base 10 logs were used)